Suppose that Steve’s marginal revenue from job search is (MR = 110 – 7.25w), where w is the wage offer at hand. Additionally, Steve’s marginal cost of job search is (MC = 60 + 2w).

a. Graph Steve’s MR and MC curves relating to job search.

b. What will Steve’s asking wage be?

c. Will Steve accept a job offer of $9.00? Justify your response with economic reasoning.

d. Suppose that unemployment insurance benefits increase, causing the marginal cost of job search to decrease to (MC = 50 + 2w). What will the new asking wage for Steve be?

e. If the replacement ratio concerning unemployment insurance were to decrease, what would be the consequences on:

i. The length of unemployment,

ii. The unemployment rate, and

iii. The post unemployment wage?

f. If the discount rate of an “individual” were to increase dramatically during the job-search process, what would be the consequences on:
i. The length of unemployment,
ii. The unemployment rate for the economy, and
iii. The post unemployment wage?
g. Define and explain the differences between frictional, seasonal, cyclical, and structural unemployment.
h. Which type of unemployment is most concerning to economists? Explain.
i. Who are the unemployed? Provide specific details using race, sex, age, and education as factors in your response.
Answer #1

Answering only first four parts as per HomeworkLib policy

A) graph

B) at equilibrium, MR & MC of job search are Equalized

So 110-7.25W = 60+2W

W = 5.405

C) now at w = 9, MR = 110-7.25*9 = 44.75

MC = 60+2w = 78

Hence MC of job search > MR from job search

So Steve will not offer the job at w = 9

D) at new equilibrium , MR = New MC

110-7.25W = 50+2W

W* = 6.486

An increase in benefits & decrease in search costs will increase the asking wage & also lengthens the duration of unemployment spell